# Prompt cusps in hierarchical dark matter halos: Implications for annihilation boost

### v1

@@ 5.3 Future directions
A definitive resolution of the order-of-magnitude difference to universal-average predictions likely requires a unified framework that embeds peak-based cusp formation criteria into merger-tree evolution with explicit environmental conditioning. There is relevant groundwork connecting peaks and excursion-set ideas, such as excursion-set peaks and related peak-constrained excursion-set formalisms [34, 33]. Complementary peak-based methods that generate halo catalogs and assembly histories directly from the initial density field—e.g. peak-patch [11] and PINOCCHIO [31, 43]—may also provide useful routes for marrying cusp-scale peak properties to hierarchical growth. Adapting such approaches to prompt cusps (including the mapping from peak parameters to cusp structural parameters and survival probabilities) is an interesting direction for future work.
@@ 6 Conclusions
  • Prompt cusps provide an additional component to the annihilation luminosity that can raise the total boost above the standard subhalo-only expectation. For fiducial survival assumptions ($f_{\rm cusp}=0.25$ and $f_{\rm surv,stripped}=1$), we find $B\sim\mathcal{O}(10)$ for Milky-Way–like hosts at $z=0$ (Fig. 2), while the no-cusp baseline remains at $B_{\rm sh}\sim\mathrm{few}$.
  • Across host masses and redshifts, prompt cusps increase the normalization of $B(M_{\rm host},z)$ while largely preserving the qualitative mass and redshift trends of the baseline prediction (Fig. 4). The dominant theoretical uncertainty is captured by the cusp-occupation fraction $f_{\rm cusp}$ and the survival of cusps associated with stripped material, $f_{\rm surv,stripped}$.
Compared to the universal-average peak-based estimate of Delos and White [18], our fiducial Milky-Way boost is lower by about one order of magnitude, which we attribute primarily to a correspondingly lower inferred number of cusps in a host. This difference highlights the importance of unifying peak-based cusp formation statistics with merger-tree evolution and environmental conditioning. Our SASHIMI-based implementation provides a flexible platform for pursuing this goal and for propagating improved cusp-formation and survival prescriptions into testable predictions for annihilation signals.
@@ Acknowledgments
This work was partly supported by MEXT KAKENHI Grant Numbers, JP20H05850, JP20H05861, and JP24K07039.

### v2

@@ 5.3 Future directions
A definitive resolution of the difference to universal-average predictions likely requires a unified framework that embeds peak-based cusp formation criteria into merger-tree evolution with explicit environmental conditioning. There is relevant groundwork connecting peaks and excursion-set ideas, such as excursion-set peaks and related peak-constrained excursion-set formalisms [37, 36]. Complementary peak-based methods that generate halo catalogs and assembly histories directly from the initial density field—e.g. peak-patch [11] and PINOCCHIO [32, 46]—may also provide useful routes for marrying cusp-scale peak properties to hierarchical growth. Adapting such approaches to prompt cusps (including the mapping from peak parameters to cusp structural parameters and survival probabilities) is an interesting direction for future work.
@@ 6 Conclusions
  • Prompt cusps provide an additional component to the annihilation luminosity that can raise the total boost above the standard subhalo-only expectation. For fiducial survival assumptions ($f_{\rm cusp}=1$ and $f_{\rm surv,stripped}=1$), we find $B\sim 50$ for Milky-Way–like hosts at $z=0$, while the no-cusp baseline remains at $B_{\rm sh}\sim\mathrm{few}$ (Fig. 2).
  • Across host masses and redshifts, prompt cusps increase the normalization of $B(M_{\rm host},z)$ while largely preserving the qualitative mass and redshift trends of the baseline prediction (Fig. 3). Our predictions are based on the fiducial assumption that prompt cusps are present in all collapsed halos and remain relevant for annihilation. The robustness of this result against modelling uncertainties is assessed by varying the cusp-occupation fraction $f_{\rm cusp}$ and the stripped-cusp contribution parameter $f_{\rm surv,stripped}$, as illustrated in Fig. 4.
Compared to the universal-average peak-based estimate of Delos and White [18], our fiducial Milky-Way boost is lower by about a factor of four, which we attribute primarily to a correspondingly lower inferred number of cusps in a host. This difference highlights the importance of unifying peak-based cusp formation statistics with merger-tree evolution and environmental conditioning. Our SASHIMI-based implementation provides a flexible platform for pursuing this goal and for propagating improved cusp-formation and survival prescriptions into testable predictions for annihilation signals.
@@ Acknowledgments
We thank M. Sten Delos for helpful discussion. This work was partly supported by MEXT KAKENHI Grant Numbers, JP20H05850, JP20H05861, and JP24K07039.